OpenStudy (anonymous):

A rectangular pen is to be built against a wall using 60 m of fencing for the three sides which need fencing. If the area of the pen is to be as large as possible, find the length of the longest side and also its maximum area.

OpenStudy (jameshorton):

450 sq.m

OpenStudy (anonymous):

I agree with @jameshorton @Devanshi

OpenStudy (anonymous):

How?

OpenStudy (jameshorton):

a = xy y + 2x = 60 y = 60 - 2x --- a = x(60 - 2x) a = 60x - 2xx --- -2xx + 60x = 0 --- the above quadratic equation is in standard form, with a=-2, b=60 and c=0 --- to solve the quadratic equation, by using the quadratic formula, copy and paste this: -2 60 0 into this solver: https://sooeet.com/math/quadratic-equation-solver.php --- the quadratic has two real roots at: --- x = 0 x = 30 --- the quadratic has a vertex maximum at: ( 15, 450 ) --- maximum area (a) occurs when x = 15: y = 60 - 2*15 y = 30 m --- answer: x = 15 m y = 30 m max area = 450 sq.m ---

OpenStudy (jameshorton):

what grade are u in

OpenStudy (anonymous):

Who me or @Devanshi

OpenStudy (jameshorton):

both of ya

OpenStudy (anonymous):

I am doing my A levels

OpenStudy (jameshorton):

oh

OpenStudy (anonymous):

and you

OpenStudy (jameshorton):

6th

OpenStudy (anonymous):

Really You are pretty smart

OpenStudy (jameshorton):

yea my mom teach me and my mom is a docter

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

How is this the standard form of equation? @lauren.ascough

OpenStudy (jameshorton):

@ganeshie8

OpenStudy (jameshorton):

@Conqueror

OpenStudy (anonymous):

So you just copy pasted the answer @jameshorton ?

OpenStudy (jameshorton):

no why

OpenStudy (jameshorton):

what

OpenStudy (anonymous):

@jameshorton did you just copy that ???

OpenStudy (jameshorton):

no

OpenStudy (anonymous):

But it is the same

OpenStudy (jameshorton):

i know

OpenStudy (anonymous):

James are you lieing

OpenStudy (jameshorton):

no

OpenStudy (anonymous):

Just explain the standard equation? @jameshorton ?

OpenStudy (anonymous):

it even has the same web address james we know you took it from there just tell us the truth

OpenStudy (anonymous):

@jameshorton come on tell us the truth

OpenStudy (jameshorton):

u want the fing truth no i didnt copy it and my real name is not jameshorton

OpenStudy (jameshorton):

i used my dad name to create that site that is my site

OpenStudy (anonymous):

lol

OpenStudy (jameshorton):

my real name is jacob

OpenStudy (jameshorton):

jacob lamb